Gross profit method

Zidek Corp. repuires an estimate of the cost of goods lost y fire on March 9. Merchandise on hand on January 1 was 38,000. Purchases since January 1 were 92.000; freight-in, 3,400; purchase returns and allowances, 2400. Sales are made at 33 1/3% above cost and totaled 120,000 to March 9. Goods costing 10,900 were left undamaged by the fire; remaining goods were destroyed.

(a) Compute the cost of goods destroyed
(b) Compute the cost of goods destroyed, assuming that the gross profit is 33 1/3 of sales.
Answer #1
a.



Merchandise on hand, January 1 

 $38.000 
Purchases

 $92.000 
Less: Purchase returns and allowances 

 $2.400 
Freight-in

 $3.400 
Total merchandise available (at cost) 

 $131.000 
Cost of goods sold

 $90.000 
Ending inventory 

 $41.000 
Less: Undamaged goods 

 $10.900 
Estimated fire loss

 $30.100 





Gross profit

25%of sales
Cost of goods sold

 $90.000 





b.



Cost of goods sold

 $79.992 
Total merchandise available (at cost)

 $51.008 
Less: Undamaged goods

 $10.900 
Estimated fire loss 

 $40.108 


answered by: Black Santa
Answer #2

SOLUTION :


a.


Stock available before fire (ending inventory (in dollars)

= (Opening stock + Purchases - Returns  and allowances + Freight-in) - COGS 

= (38000 + 92000 - 2400 + 3400) - Sales / ( 1 + 1/3)   (since sales are 33 1/3 % added to the COGS)

= 131000 - 120000 * 3/4

= 131000 - 90000

= 41000 ($)


Undamaged goods = 10900 ($)


Goods damaged in fire = 41000 - 10900 = 30100 ($) ( NSWER)


b.


CP = sales * 33 1/3 % = 120000 * 1/3 = 40000 ($)

=> COGS = Sales - GP = 120000 - 40000 = 80000


Goods available before fire 

= 131000 - 80000 

= 51000 ($)


Goods damaged în fire

= goods before fire - goods undamaged after fire

= 51000 - 10900

= 40100 ($) (ANSWER).

answered by: Tulsiram Garg